OpenStudy (anonymous):

laplace transform of cos^3at

OpenStudy (anonymous):

\[\mathscr{L}\left\{\cos^3(at)\right\}=\large\int_{0}^{\infty}e^{-st}\cos^3(at)\;dt\] The RHS looks like it'll require some tricky integration by parts. I'll see if it gets me anywhere.